OpenStudy (anonymous):

Find the product of (3x + 7y)(3x − 7y). 9x2 − 42xy + 49y2 9x2 + 42xy + 49y2 9x2 − 49y2 9x2 + 49y2

OpenStudy (anonymous):

if you know the answer could you please explain it to me? I would like to know how..

OpenStudy (anonymous):

9x2 − 49y2

OpenStudy (anonymous):

9x2 − 49y2

OpenStudy (anonymous):

(a+b)(a-b)=a^2-b^2

OpenStudy (anonymous):

(3x + 7y)(3x − 7y). First 3x times 3x = 9x2 Outside 3x times -7y = -21xy Inside 3x times 7y = 21xy Last 7y times -7y = -49y2 9 x2 - 21xy + 21xy - 49y2 = 9x2 - 49y2

OpenStudy (anonymous):

thank you!(:

OpenStudy (anonymous):

C